Please help me with my homework problem. I'm not getting the right answer. Thank you!
A chemist has 10cm3 of a 20% solution how many cubic centimeters of water should she add to produce a 5% salt solution.

Answer by ptaylor(2198)

Let x=amount of water needed
Now we know that the amount of pure salt in the 20% solution(0.20*10) plus the amount of pure salt in the water that was added(0) has to equal the amount of pure salt in the final mixture(0.05(10+x)), so our equation to solve is:
0.20*10=0.05(10+x) get rid of parens and simplify
2=0.5+0.05x subtract 0.5 from each side
2-0.5=0.5-0.5+0.05x
0.05x=1.5 divide each side by 0.05
x=30 cm^3--------------amount of water needed
CK
0.20*10=0.05*40
2=2
Hope this helps---ptaylor